# Photoluminescence of Femtosecond Laser-irradiated Silicon Carbide

### v1

@@ Abstract
Silicon carbide (SiC) is the leading wide-bandgap semiconductor material, providing mature doping and device fabrication. Additionally, SiC hosts a multitude of optically active point defects (color centers), it is an excellent material for optical resonators due to its high refractive index and an outstanding material for mechanical resonators due to its high Q/f product. Moreover, epitaxial graphene layers can be grown as ultrathin electrodes and provide the potential to fine-tune color center resonances. These characteristics render SiC an ideal platform for experiments with single color centers towards quantum technologies including coupling color centers towards cooperative effects.
A crucial step towards harnessing the full potential of the SiC platform includes technologies to create color centers with defined localization and density, e.g. to facilitate their coupling to nano-photonic structures and to observe cooperative effects. Here, silicon vacancy centers (VSi) stand out as no impurity atom is needed and high-thermal budget annealing steps can be avoided. We characterize the effect of localized, femtosecond laser irradiation of SiC, investigating surface modifications and photoluminescence including Raman spectroscopy and optical lifetime measurements.
@@ I Introduction
Silicon carbide (SiC) has seen tremendous improvements as a semiconductor platform in recent years. Together with these improvements it has also been established as a host material for single color centers and their applications in quantum technologies. Silicon vacancy centers (VSi) have several favorable characteristics to be employed in quantum technologies. In 4H-SiC, they exist in two different configurations Sörman et al. (2000) corresponding to a cubic and a hexagonal lattice site. The defects show zero phonon lines at 861 nm for V1 (cubic) and 915 nm for V2 (hexagonal) whereas the shorter wavelength falls into the detection window of standard silicon photo-detectors. Moreover, VSi do provide an optically readable ground-state spin with a zero field splitting of 4.5 MHz for V1 and 70 MHz for V2 which is straightforwardly addressable with standard RF components. However, the contrast of this optically detected magnetic resonance (ODMR) is comparably low (below 1). An additional challenge is the fact that VSi luminescence is hard to collect in (0001) SiC, because of the VSi dipole orientation in the SiC lattice. (0001) material is however the material with the highest commercial availability. Milestones include super-radiance Lukin et al. (2022) from two color centers, near-transform limited emission even in photonic nanostructures Lukin et al. (2023) as well as magnetic sensors working under extremely harsh conditions like high pressure Wang et al. (2023).
@@
Localized VSi have been created by various techniques: ion implantation (focused ion beams Wang et al. (2017) or implantation through a mask Wang et al. (2019)) or laser writing Chen et al. (2019); Castelletto et al. (2018); Zhou et al. (2023). Ion-based techniques in principle allow for the highest precision in localization which has been shown to be mainly limited by the straggle of the ions which is in the range of tens of nanometers for keV ion implantation Pezzagna et al. (2010). However, they also create high levels of damage and potential additional defects. Laser writing has been investigated as an alternative that relies on highly non-linear processes. Previous works suggest the involvement of 16 photons in the generation of a vacancy defect Chen et al. (2019). With such an extremely non-linear process, even an optical technology allows for precise localization that has been estimated to be as good as 80 nmChen et al. (2019).
@@
To address different applications in quantum technologies, not only controlling the spatial placement but also varying the density of color centers is crucial: whereas single photon sources naturally require using a single color center, higher density ensembles can be favorable for highly sensitive magnetic sensors based on optically-readable spins. In this case, higher defect density leads to brighter fluorescence and thus reduced photon shot noise in the read-out. Moreover, enhancing the density of defect ensembles may allow to observe the transition from uncoupled to collective emission (super-radiance) Gutsche et al. (2022).
@@
Laser writing can be used to create single VSi in 4H-SiC Chen et al. (2019) as well as ensembles Castelletto et al. (2018, 2020). We here perform laser processing using a commercial laser structuring system. We elucidate if laser processing might be optimized using epitaxial graphene layers that can be used as transparent electrode to electrically tune color centers. We here report on the photoluminescence from localized laser irradiated areas, we analyze surface morphology, Raman spectra as well as photoluminescence spectra and lifetimes.
@@ II Sample pre-treatment
4-Inch (0001) oriented HPSI 4H-SiC Wafers purchased from Wolfspeed were diced into quadratic chips of < 10x10 mm2. Surfaces are chemically mechanically polished to a roughness of < 0.5 nm according to manufacturer specifications. A first sample was solvent cleaned with acetone and isopropanol to remove the protective layer from dicing. It was then annealed in vacuum (5.5 x mbar) at 800°C for 3 hours. It has been observed that annealing at temperatures above 600°C reduces the initial concentration Widmann et al. (2015); Wang et al. (2019)) where ab-initio calculations show that migrate and recombine to form antisite-vacancy pairs () Bockstedte, Mattausch, and Pankratov (2004); Wang et al. (2013). On the second sample, an epitaxial graphene layer was grown at 1600°C Emtsev et al. (2009).
III Irradiation of pristine HPSI SiC
@@
We irradiate our SiC sample using an industrial-grade laser structuring system (BlueCut, Menlo Systems GmbH) with a wavelength of 1030 nm, a pulse duration of 383 fs and a repetition rate up to 1 MHz at a maximum pulse energy of 10 µJ. We focus the beam using a Thorlabs LMH-20X-1064 objective with 0.4 NA. The laser beam is linearly polarized. The focused beam had a calculated diameter of 2.4 m (). We obtain laser written patterns which we identify in our confocal laser-scanning fluorescence microscope (CLSM). FIG. 1 a) shows a white-light optical microscope image of the resulting pattern observed on the surface. The top row and the leftmost column serve as a marker, with a repetition rate of 10 kHz we used 5 pulses per spot at 1850 nJ, and 10 pulses per spot with pulse energy ranging from 60 nJ to 1850 nJ for the leftmost column to induce a clear surface modification. The rows of spots visible within this marker angle have been processed using a single pulse of the laser with pulse energies ranging from (60 - 1850) nJ. The sample was then cleaned in acetone in an ultrasonic bath to remove particles occurring during the processing. This preliminary investigation by white-light microscopy already indicates a challenge in laser writing of (dense) color center ensembles: In the regions where multiple pulses have been used for processing, around the laser processed area a structure occurs that is due to laser ablation and re-deposition of material. This re-deposited material however will not be of high quality anymore, it is typically observed to be amorphous Zhou et al. (2023).
To further analyze the structures induced on the surface especially at high pulse energies, we use atomic force microscopy (AFM) measurements carried out using a Park XE-70 AFM. The resulting image is shown in FIG. 1. b) and shows spots made with single pulses from 230 nJ to 1850 nJ. For the spots irradiated with 60 nJ, a surface modification is visible but we cannot extract its depth due to noise in our AFM. Figure 1 c) and d) summarize the depth and width of the surface structures. We add two datasets in each graph: we estimate the depth by measuring the distance between the surface and the deepest spot in the structure and we approximate the structure using a Gaussian fit. As discernible from Figure 1 c) and d) this gives us similar absolute values and a similar trend: size and width of the surface structures roughly grow with pulse energies up to 750 nJ. The diameter (width at 1/) of the features obtained above 1250 nJ with roughly 2.5 m matches the diameter of the laser beam’s Gaussian intensity profile that is 2.4 m. However, for pulse energies above 860 nJ, deviations from a Gaussian shape occur. We find a dimple with a hillock in the middle and ring shaped wall in the outer area. Due to the usage of a femtosecond laser, this is a typical mechanism when the ablation threshold is exceeded and the high energy of the pulse is able to break the covalent chemical bonds of SiC, ablating material Wang et al. (2018). In the context of localized creation of color centers, ablation processes have to be considered as a parasitic effect and will limit the creation of highly dense color center ensembles.
@@
We use a home-built confocal laser-scanning fluorescence microscopes (CLSM) to investigate the photoluminescence (PL) of the laser written color centers. For room temperature characterization, we use continuous 785 nm laser light (Coherent OBIS LX 785) for excitation. In addition, we can couple a tunable, pulsed super-continuum laser source (NKT Super K-Extreme, typical pulse width < 100 ps) which we use to measure the excited state lifetime of the observed PL. In both cases, we use a 100x microscope objective with 0.9 NA (Zeiss EC EPN 100x/0,9 DIC) to focus the laser onto the sample and to collect the PL in a reflection geometry. We separate the PL and reflected laser light using a 785 nm edge dichroic mirror (Semrock Laser-Beamsplitter HC R785 lambda/5 PV flat). We detect the PL using silicon avalanche photo-diodes (APDS) for spectrally-integrated detection (Excelitas SPCM-AQRH-14). Alternatively, we sent the PL to a spectrometer equipped with a silicon CCD (Princeton Instruments Acton Standard SP-2558 Spectrometer with Pixis 256E CCD). In both cases, we use a 850 nm long pass filter for additional laser filtering.
Figure 2 a) shows a confocal PL map of the laser processed area. Localized PL is visible resembling the regular array that was processed by irradiating the SiC sample with single laser pulses. The bottom-most row has been exposed to 60 nJ laser pulses. Here, the PL intensity is even slightly reduced compared to the surrounding SiC. This could be attributed to local annealing that can be done with low pulse energy Chen et al. (2019) and might reduce defects especially at the surface. We here compare our results to previously published results: Ref. Chen et al. (2019) identifies the process that creates in SiC as highly non-linear. According to Ref. Chen et al. (2019) 16 photons at a wavelength of 790 nm are involved in creating a single defect. For such a highly non-linear process (among other parameters), the peak intensity of the laser pulse will be crucial. We thus compare the peak intensity in our work to previous work. Ref. Chen et al. (2019) reports a minimum pulse energy for creating single to be 10.7 nJ corresponding to 22.2 TW/cm2 (at 790 nm). In Ref. Castelletto et al. (2020), the laser pulse energy necessary to create defects with a 1030 nm femtosecond laser is found to be at 58 nJ. This corresponds to 30.2 TW/cm2 and is thus slightly higher than in Ref. Chen et al. (2019). We also use 60 nJ as the lowest pulse energy. However, as we are using longer pulses and less tight focusing, in our case this only corresponds to 20.0 TW/cm2. We observe an onset of PL for the second lowest pulse energy we use which is 230 nJ corresponding to 76.6 TW/cm2. Thus, the onset of localized PL from laser irradiated areas is in accordance with previous work.

### v2

@@ Abstract
Silicon carbide (SiC) is the leading wide-bandgap semiconductor material, providing mature doping and device fabrication. Additionally, SiC hosts a multitude of optically active point defects (color centers) and is relevant for many applications in quantum technologies.
A crucial step towards harnessing the full potential of the SiC platform includes technologies to create color centers with defined localization and density, e.g. to facilitate their coupling to nano-photonic structures and to observe cooperative effects. Here, silicon vacancy centers and divacancies stand out as no impurity atom is needed and high-thermal budget annealing steps can be avoided. We characterize the effect of localized, femtosecond laser irradiation of SiC, investigating surface modifications and photoluminescence including Raman spectroscopy and optical lifetime measurements. We employ commercial high-purity, semi-insulating substrates and an industrial grade laser system to explore broader applicability of the method. As a novel approach, we apply femtosecond laser irradiation to SiC substrates with an epitaxial graphene layer and find that the threshold for photoluminescence due to laser treatment is lowered.
@@ I Introduction
Silicon carbide (SiC) has seen tremendous improvements as a semiconductor platform in recent years, especially in terms of crystal quality and wafer sizes that can now reach up to 8 inch. Together with these improvements, it has also been established as a host material for single color centers and their applications in quantum technologies. Silicon vacancy centers (VSi) have several favorable characteristics for applications in quantum technologies. In 4H-SiC, they exist in two different configurations Sörman et al. (2000) corresponding to a missing silicon atom at a cubic and a hexagonal lattice site. The defects show zero phonon lines at 861 nm for V1 (cubic) and 916 nm for V2 (hexagonal) Nagy et al. (2019); Carter et al. (2015). Moreover, VSi do provide an optically readable ground-state spin with a zero field splitting of 4.5 MHz for V1 and 70 MHz for V2 Nagy et al. (2019); Janzén et al. (2009) which is addressable with standard RF components. However, the contrast of this optically detected magnetic resonance (ODMR) is comparably low (a few ). An additional challenge is the fact that VSi luminescence is hard to collect in (0001) SiC, because of the VSi dipole orientation in the SiC lattice. (0001) material is however the material with the highest commercial availability. Milestones include super-radiance Lukin et al. (2022) from two color centers, near-transform limited emission even in photonic nanostructures Lukin et al. (2023) as well as magnetic sensors working under extremely harsh conditions like high pressure Wang et al. (2023). SiC has also emerged as a promising platform for epitaxial graphene growth by sublimation of SiC enabling the straightforward fabrication of large area, high quality graphene directly on a semiconducting substrate Berger et al. (2004); Ohta et al. (2008); Emtsev et al. (2009). Due to its high carrier mobility Novoselov et al. (2007) graphene has many possible applications in various fields Novoselov et al. (2012) such as high frequency transistors Lin et al. (2010), photonics Tamura et al. (2022); Rufangura et al. (2020) and also with VSi as transparent electrodes for Stark shift tuning of quantum systems Rühl et al. (2020).
@@
Localized VSi have been created by various techniques: ion implantation (focused ion beams Wang et al. (2017) or implantation through a mask Wang et al. (2019)) or laser writing Chen et al. (2019); Castelletto et al. (2018); Zhou et al. (2023). Ion-based techniques in principle allow for the highest precision in localization which has been shown to be mainly limited by the straggle of the ions which is in the range of tens of nanometers for keV ion implantation Pezzagna et al. (2010). However, they also create high levels of damage and potential additional defects and require costly equipment. Laser writing has been investigated as an alternative that relies on highly non-linear processes. Previous works suggest the involvement of 16 photons in the generation of a vacancy defect in SiC Chen et al. (2019). With such an extremely non-linear process, even an optical technology allows for precise localization that has been estimated to be as good as 80 nmChen et al. (2019).
@@
To address different applications in quantum technologies, not only controlling the spatial placement but also varying the density of color centers is crucial: whereas single photon sources naturally require using a single color center, higher density ensembles can be favorable for highly sensitive magnetic sensors based on optically-readable spins. In this case, higher defect density enables brighter fluorescence and thus reduces photon shot noise in the optical read-out. Moreover, enhancing defect density in ensembles may allow to observe the transition from uncoupled to collective emission (super-radiance) Gutsche et al. (2022).
@@
Laser writing can be used to create single VSi in 4H-SiC as well as ensembles Chen et al. (2019). However, the creation of dense ensembles requires higher pulse energies that are well above the ablation threshold, leading to increased damages to the crystal Wang et al. (2018). In the context of localized creation of color centers, ablation processes are a parasitic effect and will limit the creation of dense color center ensembles. We here perform laser processing using a commercial laser structuring system irradiating commercial HPSI SiC. We elucidate if laser processing is compatible with using epitaxial graphene layers that have been suggested as transparent electrode to electrically tune color centers. We here report on the photoluminescence (PL) from localized laser irradiated areas, we analyze surface morphology, Raman spectra as well as PL spectra and lifetimes.
@@ II Sample pre-treatment
A 4-inch (0001) oriented HPSI 4H-SiC wafer purchased from Wolfspeed was diced into quadratic chips of < 10x10 mm2. Diced chips were then solvent-cleaned with acetone and isopropanol to remove the protective resist layer applied for dicing. Surfaces are chemically mechanically polished to a roughness of < 0.5 nm according to manufacturer specifications. On the first sample, an epitaxial graphene layer was grown at 1600°C Emtsev et al. (2009). The second sample was annealed in vacuum (5.5 x mbar) at 800°C for 3 hours. Annealing at temperatures above 600°C has been found to reduce the concentration of Widmann et al. (2015); Wang et al. (2019) where ab initio calculations show that migrate and recombine to form pairs of antisite vacancy () Bockstedte et al. (2004); Wang et al. (2013).
III Irradiation of HPSI with graphene layer
@@
We irradiate our 4H-SiC samples using an industrial grade laser structuring system (BlueCut, Menlo Systems GmbH, repetition rate 10 kHz) with a wavelength of 1030 nm, a pulse duration of 383 fs and a repetition rate up to 1 MHz at a maximum pulse energy of 10 µJ. We focus the beam using a Thorlabs LMH-20X-1064 objective with 0.4 NA. The laser beam is linearly polarized. The focused beam had a calculated diameter of 2.4 m () in air. The sample was then cleaned in acetone in an ultrasonic bath to remove particles collected on the surface during processing. The laser-written patterns labeled pattern 1 and 2 were imaged using a commercial scanning electron microscope (SEM, Hitachi SU8000) in figure 1 a) and c), and with a white-light optical microscope in figure 1 e), corresponding to the pattern shown in figure 1 c). The top row and the leftmost column serve as markers, on pattern 1 we used 5 pulses per spot at 520 nJ on figure 1 a) (230 nJ on figure 1 c) for pattern 2), and 10 pulses per spot with pulse energy ranging from 230 nJ to 520 nJ on figure 1 a) for pattern 1 (60 nJ to 230 nJ on figure 1 c) for pattern 2) for the leftmost column to induce a clear surface modification. The rows of spots visible within this marker angle were processed using a single pulse of the laser with pulse energies ranging from 230 nJ to 520 nJ on figure 1 a) for pattern 1 (60 nJ to 230 nJ on figure 1 c) for pattern 2). This preliminary investigation by SEM and white-light microscopy already indicates a challenge in laser writing of (dense) color center ensembles: In the regions where multiple pulses have been used for processing, around the laser processed area a structure occurs that is due to laser ablation and re-deposition of material. This re-deposited material however will not be of high quality and is typically observed to be amorphous Zhou et al. (2023).
We assume that light gray areas in figure 1 a) and c) are graphene while darker parts are attributed to be the bare 4H-SiC surface. In darker areas, graphene on the surface appears to be removed, likely occurring during laser processing when ablated material was ejected scratching the surface. Additional damage may have occurred during sample handling. Although laser written spots exhibit a contrast similar to that of the surrounding graphene layer, this does not necessarily indicate that graphene remains at these sites, as all pulse energies used are well above the ablation threshold of grapheneDong et al. (2016); Roberts et al. (2011). Instead, their higher brightness compared to the bare substrate may arise from surface modification/ablation of 4H-SiC induced by laser pulses, as observed in femtosecond laser ablation of 4H-SiC studies Chen et al. (2022); Wang et al. (2022). To further study surface modifications on the graphene and 4H-SiC, atomic force microscope (AFM) measurements were performed on both patterns 1 and 2 using a commercial AFM (Park AFM XE-70). AFM images are shown in figures 1 b) and d) which are associated to laser written patterns 1 and 2 in 1 a) and c). The femtosecond laser pulses created craters on the 4H-SiC surface as observed in literature Castelletto et al. (2018). On figure 2 a) and b), we measured the depth and width (FWHM) of the craters created during the laser processing. We observe different trends for the the higher pulse-energy pattern (pattern 1) and lower pulse-energy pattern (pattern 2). On the lower energy pattern, depth and width increase up to 140 nJ and decrease until 200 nJ for the depth and 230 nJ for the width. At 230 nJ, the average depth is 40 nm bigger in the higher pulse energy pattern and the spots are 0.5 m wider. On the other hand, in the higher energy pattern, spot depth remains relatively constant with increasing pulse energy, while spots width tends to decrease with pulse energy until 480 nJ. We assume that these differing evolutions of spot dimensions as a function of pulse energy between the two patterns arise from variations in laser focusing while moving the sample to different writing locations. Figure 2 c) shows Raman spectra taken in a home-built confocal laser-scanning fluorescence microscope (CLSM), labeled CLSM 1. In CLSM 1, the sample was cooled down to 4.3 K in a closed-cycle cryostat (CryoVac). In CLSM 1, we excite the SiC sample using a tunable titanium sapphire laser (Sirah Matisse) operating at 737.19 nm focused on the sample with a microscope objective (Olympus MPLN 100x, NA = 0.9). PL and Raman scattered light is collected using the same objective and reflected laser light is rejected using a 740 nm dichroic mirror. Transmitted PL is further filtered with a longpass filter and coupled to a single mode fiber that can either be connected to an APD (Excelitas SPCM-AQRH) or a spectrometer (Horiba iHR550) equipped with a 600 lines per mm grating and a Peltier cooled CCD camera. Measurements were performed on an unprocessed part of the sample, a written spot at 520 nJ and a marker made with a train of 10 pulses at 520 nJ. The Raman lines observed on the substrate are characteristic of 4H-SiC. At 520 nJ pulse energy, the written spots exhibit Raman lines with intensities comparable to the labeled peaks from the unprocessed substrate. Whereas, compared to the substrate, TO(), LO() and TO() peaks from the marker have smaller intensities and are broadened, TO() is shifted and LA() is not visible, indicating severe damages caused by the train of 10 pulses. While this Raman spectrum shows no sign of amorphous material even for 1850 nJ, the strong decrease in Raman peak intensity has been associated with significant crystal damage in previous work Castelletto et al. (2018). The Raman spectrum does not show clear evidence for remaining graphene in accordance with the observation of ablation of material at the surface.
@@
To investigate photoluminescence from the laser processed area, we use another home-built CLSM labeled CLSM 2. In CSLM 2, we perform room temperature characterization. We use continuous 785 nm laser light (Coherent OBIS LX 785) for excitation. In addition, we can couple a tunable, pulsed super-continuum laser source (NKT Super K-Extreme, typical pulse width < 100 ps) which we use to measure the excited state lifetime of the observed PL. In both cases, we use a 100x microscope objective with 0.9 NA (Zeiss EC EPN 100x/0,9 DIC) to focus the laser onto the sample and to collect PL in a reflection geometry. We separate PL and reflected laser light using a 785 nm edge dichroic mirror (Semrock Laser-Beamsplitter HC R785 lambda/5 PV flat). We detect PL using silicon avalanche photo-diodes (APDs) for spectrally-integrated detection (Excelitas SPCM-AQRH-14). Alternatively, we sent the PL to a spectrometer equipped with a silicon CCD (Princeton Instruments Acton Standard SP-2558 Spectrometer with Pixis 256E CCD). In both cases, we use a 850 nm long pass filter for additional laser filtering. Figure 1 f) shows a room temperature confocal PL map of pattern 2 from figure 1. c). Localized PL is visible resembling the regular array that was processed by irradiating the SiC sample with single laser pulses, however, a significant level of PL also originates from areas not irradiated by the laser due to already present native defects remaining in our substrate, the PL ratio between the written spots at 60 nJ (bottom-most row) with the background is only ~2.
IV Irradiation of pristine HPSI SiC
As a reference, we treated a pristine HPSI 4H-SiC substrate using the same method described in the previous section, except using for this sample a repetition rate of 100 kHz to make the markers: we wrote a pattern on the surface with pulse energies ranging from 60 nJ (bottom row) to 1850 nJ (top row). Figure 3 a) shows a white-light optical microscope image of the resulting pattern. Figure 3 b) shows a confocal PL map of the laser processed area. The bottom-most row has been exposed to 60 nJ laser pulses, here, PL intensity is even slightly reduced compared to the surrounding, unprocessed SiC. This could be attributed to local annealing Wang et al. (2024) and surface modification as our 60 nJ pulses have energies above the reported modification threshold Chen et al. (2022) but below the ablation threshold Wang et al. (2022) necessary to create vacancies.
| Reference | Laser Focusing Objective | Comments | ||||
| Chen et al. (2019) | 790 nm | 250 fs | 1.4 NA 60x oil | 10.7 nJ | 22.2 TW/cm2 | Threshold for creating single |
| Castelletto et al. (2020) | 1030 nm | 230 fs | 0.9 NA 100x | 58 nJ | 30.2 TW/cm2 | Threshold for PL creation in the treated area |
| This work | 1030 nm | 383 fs | 0.4 NA 20x | 60 nJ | 5 TW/cm2 | Lowest pulse energy used, onset of PL for 4H-SiC with graphene |
| This work | 1030 nm | 383 fs | 0.4 NA 20x | 230 nJ | 19.2 TW/cm2 | Onset of PL for pristine 4H-SiC |
Ref. Chen et al. (2019) identifies the process that creates in SiC as highly non-linear. According to Ref. Chen et al. (2019), 16 photons at a wavelength of 790 nm are involved in creating a single defect. For such a highly non-linear process (among other parameters), the peak intensity of the laser pulse will be crucial. We thus compare the laser and optics parameters in our work to previous work in table 1. We find a slightly lower threshold than previously reported for the onset of PL on the pristine sample and a strongly reduced threshold for the sample with graphene. This could be explained by the absorptance of graphene which is 2.3 % Nair et al. (2008) compared to only 0.16% at its surface in 4H-SiC at 1030 nmYan et al. (2024) making it 14 times higher.
Figure 4. c) shows Raman spectra taken using a commercial Raman spectrometer with a 532 nm excitation laser on the pristine sample from an unprocessed area and from spots written at 60 nJ, 230 nJ and 1850 nJ. Raman spectra taken from dimples made with single pulse at 60 nJ show reduced peak intensities most notably on the TO() and LO() peaks compared to unprocessed 4H-SiC. At 230 nJ and 1850 nJ, the decrease in peak intensities is even more significant, being maximal at 1850 nJ, again indicating significant damage to the crystal lattice. To further analyze the structures, we again employ AFM measurements. The resulting image in figure 3. d) shows spots made with single pulses from 230 nJ to 1850 nJ. For the spots irradiated with 60 nJ, a shallow surface modification is visible. For spots written with pulse energy between 230 nJ to 860 nJ ablated material results in a crater-like hole. At higher pulse energy, we find a dimple with a hillock in the middle and ring shaped wall in the outer area. Figure 4 a) and b) summarize the depth and width of the surface structures. We add two datasets in each graph: we estimate the depth by measuring the distance between the surface and the deepest spot in the structure and we approximate the structure using a Gaussian fit. As discernible from figure 4 a) and b) this gives us similar absolute values and a similar trend: size and width of the surface structures roughly grow with pulse energies up to 750 nJ. The diameter (width at 1/) of the features obtained above 1250 nJ with roughly 2.5 m matches the diameter of the laser beam’s Gaussian intensity profile that is 2.4 m. However, for pulse energies above 860 nJ, deviations from a Gaussian shape occur.